OpenStudy (anonymous):

use compatible numbers to do this mentally! (+0.43) + (-1.8) + (+0.57) + (-0.2)

OpenStudy (anonymous):

hows it 0

OpenStudy (anonymous):

ok but what are the compatible numbers

OpenStudy (anonymous):

haha whoops! -1! i'm sorry, -1.8 + (-0.2) = -2

OpenStudy (anonymous):

it says make them compatible numbers first

OpenStudy (anonymous):

the answer is -1. -1.8 + (-0.2) = -2 and 0.43 + 0.57 = 1, so -2 + 1 = -1. compatible numbers means numbers that fit together, which is why i arranged the equations the way i did.